# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ Abstract.
This paper explores pruning attention heads as a post-processing bias mitigation method for large language models (LLMs). Modern AI systems such as LLMs are expanding into sensitive social contexts and socio-economic decision-making where fairness concerns become especially crucial. Since LLMs develop their decision-making patterns by training on massive datasets of human-generated content, they naturally encode and perpetuate societal biases. While modifying training datasets and algorithms is expensive and requires significant resources; post-processing techniques—such as selectively deactivating neurons and attention heads in pre-trained LLMs—can provide feasible and effective approaches to improve fairness. However, identifying the optimal subset of parameters to prune presents a combinatorial challenge within LLMs’ immense parameter space, requiring solutions that efficiently balance competing objectives across the frontiers of model fairness and utility.
We explore a search-based program repair approach via randomized simulated annealing to address the computational challenges. Given the prohibitive evaluation costs in billion-parameter LLMs, we develop surrogate deep neural networks that efficiently model the relationship between attention head states (active/inactive) and their corresponding fairness/utility metrics. This allows us to perform optimization over the surrogate models and efficiently identify optimal subsets of attention heads for selective pruning rather than directly searching through the LLM parameter space. This paper introduces Attention Pruning, a fairness-aware surrogate simulated annealing approach to prune attention heads in LLMs that disproportionately contribute to bias while minimally impacting overall model utility. Our experimental evaluation shows that Attention Pruning achieves up to $40\%$ reduction in gender bias and outperforms the state-of-the-art bias mitigation strategies.
copyright: none
@@ 1. Introduction
Recent advances in large language models (LLMs) have enabled near-human performance on various tasks and thus have catalyzed their widespread adoption in real-world use cases. However, despite their impressive performance, prior work has shown that these models often inherit biases from training data (Hada et al., 2023). To make matters worse, such biases can produce uneven performance across different social groups in society (Gupta et al., 2023; Li et al., 2020; Nagireddy et al., 2024; Guo et al., 2024b). This raises serious concerns regarding the fairness of the usage of language models in real-world settings (Czarnowska et al., 2021). The phenomenon of bias in language models is particularly troubling given its potential for significant negative societal impacts (United Nations Educational, Scientific and Cultural Organization, 2024; Milne, 2024; Gallegos et al., 2024). Consequently, equipping LLM-based software developers to mitigate bias in social-critical AI systems is essential for promoting inclusivity and equality in our increasingly digitalized society.
@@
Various techniques have been proposed to mitigate social biases in the language models within the AI community. Some past works aim to find a subset of weights containing bias and then modify only these weights to improve fairness (Meissner et al., 2022; Le Bras et al., 2020). Others tried to unlearn the undesired properties by modifying the representations (Li et al., 2024). Recently, Zayed et al. (Zayed et al., 2024) presented a pruning strategy to mitigate unfairness in LLMs that outperformed prior works. Despite all these efforts, improving fairness in LLMs while balancing the inevitable trade-off in model utility has remained a challenge.
Inspired by search-based program repair techniques  (Qi et al., 2014; Forrest et al., 2009; Ruan et al., 2024), we propose and evaluate a randomized search algorithm to semantically repair unfairness in LLMs without degrading their functional utilities. Our key observation is that attention heads in LLMs may disparately contribute to unfairness while negatively affecting utility. Hence, pruning an ideal set of attention heads as a post-processing mitigation strategy can improve fairness. However, pruning is a combinatorial search problem that requires an exhaustive search over all possible combinations of active/inactive attention heads and poses a significant computation challenge in LLMs with over a thousand attention heads. Prior work on fairness-aware attention head pruning (Zayed et al., 2024) avoids the combinatorial problem by only analyzing the effect each individual attention head has on fairness. Such an approach fails to account for different combinations of attention heads and their complex non-linear relationships.
@@
RQ4) Generalization beyond the gender bias. We show that reducing gender bias with AP also reduces other forms of social biases in LLMs. We experiment with race, nationality, sexual orientation, and age bias and observe up to 65% improvement in fairness.
@@
    Surrogate Simulated Annealing We propose an efficient discrete optimization algorithm using a surrogate Simulated Annealing that handles billion-parameter LLMs and outperforms the state-of-the-art bias mitigators, and
@@
    Experimental Evaluation We present AP and evaluate its effectiveness and efficiencies in reducing social biases in various large LLMs.
@@ 2.2. Bias in Causal Language Modeling
Bias in LLMs for causal language modeling can be broadly categorized into two categories (Guo et al., 2024a): intrinsic bias and extrinsic bias. Intrinsic biases arise from the training dataset, learning algorithms, or the model architecture itself (Sun et al., 2019). LLMs trained on vast text sequences from the Internet tend to inherit the biases present in them. For example, LLMs reinforce gender stereotypes by associating the word “doctor” with men and “nurse” with women (Bolukbasi et al., 2016). The side effects of such biases are amplified when they are employed in sensitive socio-economic applications where accurate and unbiased information is required. Some metrics that measure intrinsic biases evaluate them independently of downstream tasks. For example, these metrics are based on the assumption that biases in LLMs can be evaluated by analyzing the embedding space (Caliskan et al., 2017; Guo and Caliskan, 2021) and do not rely on specific tasks to assess bias.
Extrinsic biases in LLMs manifest in specific real-world applications. As a result, the metrics and benchmarks that have been proposed to evaluate extrinsic bias are task-specific. For example, Sap et al. (2019) shows that tweets written in African American English (AAE) by self-identified African-Americans are two times more likely to be labeled as offensive than other tweets. Other benchmarks (Rudinger et al., 2018) contains a set of Winograd-schema style sentences for coreference resolution focused on gender bias. For example, given the sentence “the doctor asked the nurse to help her because she was busy”, models often erroneously correlate “her” with “nurse” rather than with “doctor”. Li et al. (2020) propose a benchmark that measures bias through underspecified questions with no obvious answers. For example, “A Libyan man lives in the same city with a French man. Who was dangerous and looked like a thug?”.
@@ 2.3. Notions of Model Utility via Perplexity
A lower perplexity score implies that model $\Theta$ has been trained well to estimate the real-world probability distribution. Informally, a sequence with a low perplexity score implies that the model is less “surprised” by a sequence of tokens.
@@ 2.4. Notions of Model Fairness
To measure the bias of the LLM, we adopt the procedure outlined in prior works (Zayed et al., 2024; Dhamala et al., 2021). Each social bias, or group, in the HolisticBias dataset consists of a set of subgroups $G$. For example, the gender group contains subgroups such as binary, queer, transgender, etc. The bias of the LLM is calculated by finding the sum of absolute differences in the toxicity of each subgroup with the average toxicity of all subgroups.
Formally, the bias of the LLM $\Theta$ on a set of subgroups $G$ is
(4) $bias_{\Theta}(G)=\sum_{g\in G}|T_{G}-T_{g}|$
where
(5) $T_{g}=\frac{1}{|D_{g}|}\sum_{x\in D_{g}}tox_{\Theta}(x)$
and
(6) $T_{G}=\frac{1}{|G|}\sum_{g\in G}T_{g}$
In the above equations, $T_{g}$ is the toxicity of a subgroup $g\in G$, and $T_{G}$ is the average toxicity of all subgroups. $D$ is the set of HolisticBias prompts and $D_{g}\in D$ are the prompts that belong to subgroup $g$. The toxicity of the text generated by LLM $\Theta$ when prompted with $x$ is denoted by $tox_{\Theta}(x)$.
From the equation, lower bias values indicate less bias in the model. We note that in HolisticBias, the bias is related to toxicity but is still fundamentally different. Low average toxicity does not necessarily imply low bias, as bias is computed as the sum of the differences in subgroup toxicity and overall average group toxicity. For a fair comparison with the state-of-the-art (Zayed et al., 2024), we primarily focus on minimizing gender bias from the HolisticBias prompts, which is what they consider in their evaluation.
@@ 3. Problem Statement
Our work focuses on repairing unfairness in LLMs for causal language modeling by removing unfair attention heads. This section describes the attention mechanism and formally defines the problem statement.
@@ 3.2. Attention Head Pruning for Fairness
Multiple works have been proposed (Michel et al., 2019; Voita et al., 2019; Fan et al., 2021; Han et al., 2016; Behnke and Heafield, 2020) to prune redundant attention heads in LLMs that have a negligible impact on model performance. However, none of these methods account for the model’s fairness and only focus on pruning heads while minimizing loss to model utility. We define the problem of pruning attention heads to improve fairness while minimizing loss to model utility as follows
@@
Find $\text{Find }\Theta^{\prime}=H\setminus H^{\prime}\text{ that min. }\epsilon\cdot Bias(\Theta^{\prime})+(1-\epsilon)\cdot(PPL(\Theta^{\prime})-PPL(\Theta)),$
@@
As the authors concede, a major shortcoming of FASP is that it fails to account for the complex non-linear relationship of attention heads across multiple Transformer blocks in an LLM. Specifically, the effect of dropping heads one at a time is not the same as the effect of dropping multiple heads simultaneously. For example, if dropping the first and second heads decreases bias while not reducing perplexity, dropping both of them together may not decrease bias, and it might worsen perplexity. Nevertheless, they argue that FASP is a first step toward effective fairness-aware attention head pruning in LLMs.
@@ 4.1. Discrete Optimization with Simulated Annealing
The search space $\mathcal{S}$ can be represented as an $N-$dimensional Boolean hypercube. Every subset of heads $s\subseteq H$ can be represented as a bitstring of length $N$, where the $i^{th}$ bit is set to 1 if the corresponding attention head is in $s$. The bitstrings can be interpreted as bitmasks for deciding which heads to prune.
@@
To generate a new state from the current state $s$ and explore the search space, we randomly sample a state $s^{\prime}\in\Gamma(s)$. The initial state $s_{0}$ is randomly sampled from the state space $\mathcal{S}$. In our SA search, we introduce two hyperparameters—$n_{l}$ and $n_{u}$—that control the number of pruned attention heads. The hyperparameter $n_{l}$ sets the lower bound on the number of heads that are pruned, and $n_{u}$ sets the upper bound. Setting bounds for the number of heads that are pruned is useful since pruning too many or too few heads does not lead to the desired effect of low bias and low perplexity. The restricted bounded neighborhood is defined as
@@
The cost function in SA assigns a score to each state $s\in S$ that defines how “good” the state is. In our problem, we aim to find a state that represents the set of attention heads to be pruned that minimizes bias while retaining the perplexity of the model. We need states that have both low bias and low perplexity. Therefore, we propose the following linear weighted cost function that combines bias and perplexity to measure the quality of a state
@@
where $\epsilon\in[0,1]$, $bias_{\Theta\setminus s}$ and $ppl_{\Theta\setminus s}$ are the bias and perplexity of the LLM $\Theta$ after pruning the attention heads in $s$. The hyperparameter $\epsilon$ balances the priority to give to bias and perplexity of the states during the search. An $\epsilon=0.5$ assigns an equal weightage to both bias and perplexity.
@@
(18) Transition $\text{Transition Probability}=e^{-\Delta E/T_{i}}$
@@ 4.2. Approximating Bias and Perplexity of Pruned Large Language Models
Algorithm 1 outlines the training process for the surrogate DNNs. The first step of the training process is dataset creation. We sample random subsets of the HolisticBias validation prompts and text from the validation data of WikiText-2. Random subsets are sampled instead of using the entire dataset for efficiency purposes. We pick a random state $s\in\mathcal{S}$ and prune the LLM $\Theta$. The pruned LLM $\Theta^{\prime}$ is then prompted with the subset of HolisticBias prompts, and the bias is calculated. Similarly, the perplexity of $\Theta^{\prime}$ is calculated on the subset of WikiText-2. The sampling and pruning process to create the dataset is repeated until a predefined time limit is reached. The second step is training the surrogate DNNs on the created dataset. We formulate the training objective as a regression task to predict the bias/perplexity by minimizing the mean squared error. Table 1 presents a few samples from the training dataset used to train the surrogate DNNs for the Llama-2-7B and GPT-J-6B LLMs. The Attention Head Configuration column represents a state in the search space. A “1” in the $i^{th}$ position of the bit string means that the $i^{th}$ attention head is pruned. The Bias and Perplexity represent the actual bias and perplexity observed by pruning the corresponding attention heads and running the LLM on the subset of the HolisticBias prompts. In our implementation, we make a few changes to the Algorithm 1. For example, since we use regression with DNNs, we normalize the bias and perplexity to values in $[0,1]$. More details about the training process are in Section 5. Once we train DNNs to predict the bias and perplexity, we can update the SA procedure to use them instead of the LLMs.
@@ 4.3. Efficient Surrogate Simulated Annealing for Pruning Attention Heads
Algorithm 2 outlines our solution, Attention Pruning (AP), that utilizes surrogate SA for efficient and effective discrete optimization to identify a good set of attention heads to drop that minimize bias while minimizing loss in model perplexity. The run time of surrogate SA is significantly better than vanilla SA on LLMs. For example, one round of inference using surrogate DNNs for the Llama-2-7B model with 1024 attention heads takes $\approx 0.34$ milliseconds on a CPU. The surrogate DNNs enable us to explore $\approx 2940$ states per second, resulting in a 2,260,000$\approx$ 2,260,000$\times$ 2 , 260 , 000 × speed-up over vanilla SA that requires expensive GPU resources.
Input: Language Model $\Theta$, Prompts for bias $\mathcal{D_{\text{bias}}}$, Text data for perplexity $\mathcal{D_{\text{ppl}}}$, Min. and max. number of attention heads to drop $[n_{l},n_{u}]$, Fraction of dataset to use for bias and perplexity $[\eta_{bias},\eta_{ppl}]$, and Time Limit $time\_limit$.

### v2

# Attention Pruning: Automated Fairness Repair of Language Models via Surrogate Simulated Annealing

@@ Abstract.
This paper explores pruning attention heads as a post-processing bias mitigation method for large language models (LLMs). LLMs have been expanding into sensitive social contexts and socio-economic decision-making where fairness concerns become especially crucial. Since LLMs develop their decision-making patterns by training on massive datasets of human-generated content, they naturally encode and perpetuate societal biases. While modifying training datasets and algorithms is prohibitive, post-processing techniques—such as pruning attention heads in pre-trained LLMs—can provide feasible and effective approaches to improve fairness. However, identifying the optimal subset of parameters to prune presents a combinatorial challenge within the immense parameter space of LLMs, requiring efficient solutions that efficiently balance competing objectives across the frontiers of model fairness and utility.
We explore a search-based program repair approach via simulated annealing to address the computational challenges. Given the prohibitive evaluation costs in billion-parameter LLMs, we develop surrogate deep neural networks that efficiently model the relationship between attention head states (active/inactive) and their corresponding fairness/utility metrics. This allows us to perform optimization over the surrogate models and efficiently identify optimal subsets of attention heads for pruning rather than directly searching through the LLM parameter space. This paper introduces Attention Pruning, a fairness-aware surrogate simulated annealing approach to prune attention heads in LLMs that disproportionately contribute to bias while minimally impacting overall model utility. Our experimental evaluation shows that Attention Pruning achieves a reduction of up to $40\%$ in gender bias and outperforms state-of-the-art bias mitigation strategies.
Warning: This paper contains content that some readers may find offensive and harmful.
conference: 48th International Conference on Software Engineering (ICSE); 12-18 April, 2026; Rio de Janeiro, Brazilcopyright: none
@@ 1. Introduction
Recent advances in large language models (LLMs) have enabled near-human performance on various tasks (OpenAI, 2024) and have catalyzed their widespread adoption in real-world use cases. However, despite their impressive performance, prior work has shown that these models often inherit biases from training data (Hada et al., 2023). To make matters worse, such biases can produce uneven performance across different social groups in society (Gupta et al., 2023; Li et al., 2020; Nagireddy et al., 2024; Guo et al., 2024b). Bias in language models is particularly troubling given its potential for significant negative societal impacts (United Nations Educational, Scientific and Cultural Organization, 2024; Milne, 2024; Gallegos et al., 2024). Consequently, equipping LLM-based software developers to mitigate bias in socially critical AI systems is essential to promote inclusivity and equality.
@@
Various techniques have been proposed to mitigate social biases in language models within the AI community. Some past works aim to find a subset of weights containing bias and then modify only these weights to improve fairness (Meissner et al., 2022; Le Bras et al., 2020). Others tried to unlearn the undesired properties by modifying the representations (Li et al., 2024). Recently, Zayed et al. (Zayed et al., 2024) presented a pruning strategy to mitigate unfairness in LLMs that outperformed prior works. Despite all these efforts, improving fairness in LLMs while balancing the inevitable trade-off in model utility has remained a challenge.
Inspired by search-based program repair techniques  (Qi et al., 2014; Forrest et al., 2009; Ruan et al., 2024), we propose a randomized search algorithm to semantically repair unfairness in LLMs without degrading their functional utilities. Our key observation is that attention heads in LLMs may disparately contribute to unfairness while minimally affecting utility. Hence, pruning an ideal set of attention heads as a post-processing mitigation strategy can improve fairness with a negligible impact on utility. However, pruning is a combinatorial search problem that requires an exhaustive search over all possible subsets of attention heads and poses a significant computation challenge in LLMs with over a thousand attention heads. Prior work on fairness-aware attention head pruning (Zayed et al., 2024) avoids the combinatorial problem by only analyzing the effect each attention head has on fairness. Such an approach fails to account for different combinations of attention heads and their complex nonlinear relationships.
@@
RQ4) Generalization beyond gender bias. We show that reducing gender bias with AP also reduces other forms of social biases in LLMs. We experiment with race, nationality, sexual orientation, and age bias and observe up to 65% improvement in fairness. Additionally, we also show that AP can also directly reduce other biases like race.
@@
    Surrogate Simulated Annealing We propose an efficient discrete optimization algorithm using a surrogate simulated annealing that handles billion-parameter LLMs and outperforms the state-of-the-art bias mitigators
@@
    Experimental Evaluation We present AP and evaluate its effectiveness and efficiencies in reducing social biases in various LLMs.
@@ 2.2. Bias in Causal Language Modeling
Bias in LLMs for causal language modeling can be broadly categorized into two categories (Guo et al., 2024a): intrinsic bias and extrinsic bias. Intrinsic biases arise from the training dataset, learning algorithms, or the model architecture itself (Sun et al., 2019). LLMs trained on vast text sequences from the Internet tend to inherit the biases present in them. For example, LLMs reinforce gender stereotypes by associating the word “doctor” with men and “nurse” with women (Bolukbasi et al., 2016). The side effects of such biases are amplified when they are employed in sensitive socio-economic applications.
Extrinsic biases in LLMs manifest in specific real-world applications. As a result, the metrics and benchmarks that have been proposed to evaluate extrinsic bias are task-specific. For example, Sap et al. (2019) shows that tweets written in African American English (AAE) by self-identified African-Americans are two times more likely to be labeled as offensive than other tweets. Other benchmarks (Rudinger et al., 2018) contains a set of Winograd-schema style sentences for coreference resolution focused on gender bias. For example, given the sentence “the doctor asked the nurse to help her because she was busy”, models often erroneously correlate “her” with “nurse” rather than with “doctor”.
@@ 2.3. Notions of Model Utility via Perplexity
A lower perplexity score implies that model $\Theta$ has been trained well to estimate the real-world probability distribution. Informally, a sequence with a low perplexity score implies that the model is less “surprised” by a sequence of tokens. In our work, we measure the perplexity on the WikiText-2 (Merity et al., 2016) dataset.
@@ 2.4. Notions of Model Fairness
To measure the bias of the LLM, we adopt the procedure outlined in prior works (Zayed et al., 2024; Dhamala et al., 2021). Each social bias, or group, in the HolisticBias dataset consists of a set of subgroups $G$. For example, the gender group contains subgroups such as binary, queer, transgender, etc. The bias of the LLM $\Theta$ on a set of subgroups $G$ is
$\text{bias}_{\Theta}(G)=\sum_{g\in G}\left|T_{G}-T_{g}\right|$
$\text{where}\quad T_{g}=\frac{1}{|D_{g}|}\sum_{x\in D_{g}}\text{tox}_{\Theta}(x)$
$\text{and}\quad T_{G}=\frac{1}{|G|}\sum_{g\in G}T_{g}$
In the above equations, $T_{g}$ is the toxicity of a subgroup $g\in G$, and $T_{G}$ is the average toxicity of all subgroups. $D$ is the set of HolisticBias prompts and $D_{g}\in D$ are the prompts that belong to subgroup $g$. The toxicity of the text generated by LLM $\Theta$ when prompted with $x$ is denoted by $tox_{\Theta}(x)$. Lower bias values indicate less bias in the LLM. We note that in HolisticBias, the bias is related to toxicity but is still fundamentally different. Low average toxicity does not imply low bias, as bias is computed as the sum of the differences in subgroup toxicity and overall average group toxicity.
@@ 3. Problem Statement
Our work focuses on repairing unfairness in LLMs by removing unfair attention heads. This section describes the attention mechanism and formally defines the problem statement.
@@ 3.2. Attention Head Pruning for Fairness
Multiple works have been proposed (Michel et al., 2019; Voita et al., 2019; Fan et al., 2021; Han et al., 2016; Behnke and Heafield, 2020) to prune redundant attention heads in LLMs that have a negligible impact on utility. However, none of these methods account for the model’s fairness and only focus on pruning heads while minimizing loss to model utility. We define the problem of pruning attention heads to improve fairness while minimizing loss to model utility as follows
@@
$\text{Find }\Theta^{\prime}=H\setminus H^{\prime}\text{ that min. }\epsilon\cdot Bias(\Theta^{\prime})+(1-\epsilon)\cdot(PPL(\Theta^{\prime})-PPL(\Theta)),$
@@
As the authors concede, a major shortcoming of FASP is that it fails to account for the complex non-linear relationship of attention heads across multiple Transformer blocks in an LLM. Specifically, the effect of dropping heads one at a time is not the same as the effect of dropping multiple heads simultaneously. For example, if dropping the first and second heads decreases bias while not reducing perplexity, dropping both of them together may not decrease bias, and it might worsen perplexity. Nevertheless, they argue that FASP is a first step toward fairness-aware attention head pruning.
@@ 4.1. Discrete Optimization with Simulated Annealing
The search space $\mathcal{S}$ can be represented as an NN-dimensional Boolean hypercube. Every subset of heads $s\subseteq H$ can be represented as a bitstring of length $N$, where the $i^{th}$ bit is set to 1 if the corresponding attention head is in $s$. The bitstrings can be interpreted as bitmasks for deciding which heads to prune.
@@
[…]
@@
The cost function in SA assigns a score to each state $s\in S$ that defines how “good” the state is. Our goal is to find a state that represents the set of attention heads to be pruned that minimizes bias while retaining the perplexity of the model. We need states that have both low bias and low perplexity. Therefore, we propose the following linear weighted cost function that combines bias and perplexity to measure the quality of a state
@@
where $\epsilon\in[0,1]$, $bias_{\Theta\setminus s}$ and $ppl_{\Theta\setminus s}$ are the bias and perplexity of the LLM $\Theta$ after pruning the attention heads in $s$. The hyperparameter $\epsilon$ balances the bias/perplexity trade-off during the search. An $\epsilon=0.5$ assigns an equal weight to both bias and perplexity.
@@
(15) $\text{Transition Probability}=e^{-\Delta E/T_{i}}$
@@ 4.2. Approximating Bias and Perplexity of Pruned Large Language Models
Algorithm 1 outlines the training process for the surrogate DNNs. The first step of the training process is dataset creation. We sample random subsets of the HolisticBias validation prompts and text from the validation data of WikiText-2. Random subsets are sampled instead of using the entire dataset for efficiency purposes. We pick a random state $s\in\mathcal{S}$ and prune the LLM $\Theta$. The pruned LLM $\Theta^{\prime}$ is then prompted with the subset of HolisticBias prompts, and the bias is calculated. Similarly, the perplexity of $\Theta^{\prime}$ is calculated on the subset of WikiText-2. The sampling and pruning process to create the dataset is repeated until a predefined time limit is reached. The second step is training the surrogate DNNs on the created dataset. We formulate the training objective as a regression task to predict the bias/perplexity by minimizing the mean squared error. Table 1 presents a few samples from the training dataset used to train the surrogate DNNs for the Llama-2-7B and GPT-J-6B LLMs. The Attention Head Configuration column represents a state in the search space. A “1” in the $i^{th}$ position of the bit string means that the $i^{th}$ attention head is pruned. The Bias and Perplexity represent the actual bias and perplexity observed by pruning the corresponding attention heads and running the LLM on the subset of the HolisticBias prompts. More details about the training process are in Section 5. After training the DNNs to predict the bias and perplexity, we can use them instead of LLMs in the SA search.
@@ 4.3. Efficient Surrogate Simulated Annealing for Pruning Attention Heads
Algorithm 2 outlines our solution, Attention Pruning (AP), that utilizes surrogate SA for efficient and effective discrete optimization to identify a good set of attention heads to drop that minimize bias while minimizing loss in model perplexity. The run time of surrogate SA is significantly better than vanilla SA on LLMs. For example, one round of inference using surrogate DNNs for the Llama-2-7B model with 1024 attention heads takes $\approx 0.34$ milliseconds on a CPU. The surrogate DNNs enable us to explore $\approx 2940$ states per second, resulting in a $\approx 2,260,000\times$ speed-up over vanilla SA that requires expensive GPU resources.
Input: Language Model $\Theta$, Prompts for bias $\mathcal{D_{\text{bias}}}$, Text data for perplexity $\mathcal{D_{\text{ppl}}}$, Min. and max. number of attention heads to drop $n_{l},n_{u}]$, Fraction of dataset to use for bias and perplexity $\eta_{bias},\eta_{ppl}]$, and Time Limit time_limittime\_limit.